OpenStudy (anonymous):

Find an explicit and a recursive formula for each geometric sequence: 16,4,1,1/4,1/16 Explicit formula Recursive formula 5,15,45,135,405 Explicit formula Recursive formula

Parth (parthkohli):

1) \(\Large \color{MidnightBlue}{\Rightarrow 64 \div 4^n }\) \(\Large \color{MidnightBlue}{\Rightarrow 16 \div 4^{n - 1} }\) Can you tell which is explicit and which is recursive?

OpenStudy (anonymous):

Explicit: \(T_n = a_1 \times r^{n-1} \) Recursive: \(T_1= 16\) and \( \forall n >1,\in \mathbb{N} ,\;\; T_{n+1} = T_n\times r\)

OpenStudy (anonymous):

\(a_1 =\) First term \(T_n =\) nth term \(r=\) common ratio